OpenStudy (anonymous):

solve the equation for for 0° ≤ x ≤ 360° sin 2x - sin 3x + sin 4x = 0 Need help. Thanks!

OpenStudy (anonymous):

use sinA+sinB=2sin(A+B/2)sin(A-B/2)

OpenStudy (raden):

that's wrong identity ^ sinA + sinB = 2 sin((A+B)/2)cos((A-B)/2)

OpenStudy (raden):

first, simplify that sin2x + sin4x = 2sin((2x+4x)/2)cos((2x-4x)/2) = 2sin(3x)cos(-x) = 2sin3xcosx

OpenStudy (raden):

so, sin 2x - sin 3x + sin 4x = 0 2sin3xcosx - sin3x = 0 factor out! sin3x (2cosx - 1) = 0 set each factour equal zero then solve for x sin3x = 0 x = ..... 2cosx - 1 = 0 cosx = 1/2 x = .....

OpenStudy (anonymous):

oh yes..i'm sorry! my mistake.

OpenStudy (anonymous):

got it! thanks !!